July 2013 Reports

To the main

29th July 2013 (Monday)

  • Mostly MC productions ongoing, tail of reprocessing and restripping campaign
  • T0:
    • CERN: NTR
  • T1:

25th July 2013 (Thursday)

  • Mostly MC productions ongoing, tail of reprocessing and restripping campaign
  • T0:
    • CERN: Issues with Oracle DBs
  • T1:

15th July 2013 (Monday)

  • Incremental stripping campaign in progress and MC productions ongoing
  • T0:
    • CERN: Wrongly terminated FTS transfers (GGUS:95642)
  • T1:
    • GridKa: currently we have no problem in staging, but we have no any information what was changed (GGUS:95135)
    • PIC: Pilots aborted (GGUS:95730) solved by removing all *_sl5 queues from configuration.

8th July 2013 (Monday)

  • Incremental stripping campaign in progress and MC productions ongoing
  • T0:
  • T1:

3d July 2013 (Thursday)

  • Incremental stripping campaign in progress and MC productions ongoing
  • T0:
  • T1:
    • GridKa: Problem with staging was solved (GGUS:95135), still we have backlog of jobs.
    • GridKa: Pilots submitted through KIT brokers Failed(8%) or Cleared (70%) (GGUS:94462)
    • RAL: CVMFS problem (GGUS:95435).
    • CNAF: Problem with FTS transfers today. Fixed with roll back to the previous SRM release.

1st July 2013 (Monday)

  • Incremental stripping campaign in progress and MC productions ongoing
  • T0:
  • T1:
Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2013-09-02 - JoelClosier
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback